1. Use the information in the following table to calculate: a. the adult population
b. the labor force
c. the labor-force participation rate

d. the unemployment rate

Employed

Unemployed

Not in the labor force

142,263,000

10,112,000

82,932,000

2. Assume that the reserve requirement is 3 percent. All other equal, will the money supply expand more if the Federal Reserve buys $3,000 worth of bonds or if someone deposits in a bank $3,000 that he had been hiding in his cookie jar? If one creates more, how much more does it create? Explain your reasoning.

3. Assume that the reserve requirement is 10 percent. Also assume that banks do not hold excess reserves and there is no cash held by the public. The Federal Reserve decides that it wants to decrease the money supply by $20 million.

a. If the Fed is using open-market operations, will it buy or sell bonds?
b. What quantity of bonds does the Fed need to buy or sell to accomplish the goal? Explain your

reasoning.

4. Suppose that money supply is $4 trillion, nominal GDP is $20 trillion, and real GDP is $16 trillion.

a. What is the price level? What is the velocity of money?

Suppose that velocity is constant and the economy’s output of goods and services rises by 3 percent each year.

b. What will happen to nominal GDP and the price level next year if the Fed increases the money supply by 5 percent?

c. What money supply should the Fed set next year if it wants to keep the price level stable? d. What money supply should the Fed set next year if it wants inflation of 1.5 percent?

5. The following table reports the price of a McDonald’s Big Mac for six different countries, with each price denominated in units of the local currency. The table also reports the current-marketnominal exchange rate between the United States and the five remaining countries.

Country

U.S.A.
Chile
Hungary
Czech Republic Brazil

Canada

Price

$3.99 1,800 pesos 700 forints 65 korunas 8.1 real
4.0 C$

Current-market nominal exchange rate

-- 552 pesos/$

200 forints/$ 18.7 korunas/$ 2.2 real/$
1.2 C$/$

1

a. Compute the U.S. dollar nominal exchange rate that is implied by the theory of purchasing- power parity (PPP) for each country. Remember, the nominal exchange rate is expressed as units of foreign currency per U.S. dollar.

b. For each country, compute the actual real exchange rate measured as the number of foreign Big Mac per U.S. Big Mac.

c. According to PPP, what is the predicted nominal exchange rate between the koruna and the Canadian dollar (i.e., koruna per C$)? Based on the data given in the table, what is the actualcurrent-market nominal exchange rate between these two currencies?
Answer #1

1.

a. The adult population = employed + unemployed + not in the labor force = 142,263,000 + 10,112,000 + 82,932,000
= 235,307,000

b. The labor force = Employed + Unemployed = 142,263,000 + 10,112,000 = 152,375,000

c. The labor-force participation rate = (Labor force/Adult population)*100 = (152,375,000/235,307,000)*100 = 64.76%

d. The unemployment rate = (Unemployed/Labor force)*100 = (10,112,000/152,375,000)*100 = 6.64%

(Note: Post 1 question at a time.)